OpenStudy (anonymous):

Find the probability for the experiment of drawing two marbles (without replacement) from a bag containing four green, four yellow, and four red marbles. Both marbles are green.

OpenStudy (anonymous):

You start off with 12 marbles off which 4, i.e 1/3, are green

OpenStudy (anonymous):

so for the first marble, you have a 1/3 chance of getting a green

OpenStudy (anonymous):

now we have 12-1=11marbles left in the bag, of which 3 are green, so the chance of getting a green marble this time is 3/11

OpenStudy (anonymous):

The probability of getting both of these correct is the product of their individual products,, i.e (3/11)*(1/3)

OpenStudy (anonymous):

can you find what that is equal to?

OpenStudy (anonymous):

\[\frac{ 3 }{ 11 } \times\frac{ 1 }{ 3 } = \frac{ 3\times1}{3\times11 } = \frac{ 1 }{ 11 }\]